OpenStudy (anonymous):

A quarry shaped like a cone has a height of 42 ft and a diameter of 48 ft. What is the volume of the quarry? Use 3.14 to approximate pi, and express your final answer in hundredths. ft^3

OpenStudy (anonymous):

@StudyGurl14 @Data_LG2

jimthompson5910 (jim_thompson5910):

Volume of cone \[\Large V = \frac{1}{3}*\pi*r^2*h\] V = volume pi = 3.14 approximately r = radius h = height

jimthompson5910 (jim_thompson5910):

radius = diameter/2

OpenStudy (anonymous):

@jim_thompson5910 what would be the radius?

jimthompson5910 (jim_thompson5910):

you take the diameter and cut it in half to get the radius

OpenStudy (anonymous):

ok so its 24

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

so 3.14*24^2*42

OpenStudy (anonymous):

75962.88

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Data_LG2

jimthompson5910 (jim_thompson5910):

I'm getting the same

OpenStudy (anonymous):

ok can you help wih 3 more

jimthompson5910 (jim_thompson5910):

I'll help with one more

jimthompson5910 (jim_thompson5910):

which of the 3 is giving you the most trouble?

OpenStudy (anonymous):

i will get it i will tag you in it

jimthompson5910 (jim_thompson5910):

ok